• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Rvachev

    Log entry time 07:56:26 on October 26,1999

    Entry number 29215

    keyword=Shift summary

    12:01 beam off for maintenance

    12:03 target masked and moved to empty position; the fan set to 6%,
    JT valve closed a bit

    12:03-8:00 coolant flow to hall A oscillates between ~ 6 and 7.5 g/s
    with period ~ 2min, sometimes oscillations sieze; the oscillations seem
    to be correlated with ESR CPI673 pressure sensor readings; coolant temperature
    stable at 13.5K (at ESR).

    6:50 Hall A went to controlled access

    7:20 Hall A went to restricted access